Using a calculator, find log 10000 where log means log to the base of ten.Show work

Answer by funmath(2925)

There seams to be a conflict in your directions. "Use your calculator" and "show your work"
10000=10^4
log 10000=log10^4
Log (base a)(a^y)=y so
log (base 10)(10^4)=4
Put it in your calculator to check.
log(10000)=4